A school record in a certain race in 1960 was 3.8 minutes. In 1980 the school record in the same event was 3.65 minutes.Choose the linear equation for these data points.(y=-.0075x+18.5),(y=.0075x-18.5), (y=.0075x 0.18.5)
Answer by Boreal(15235)

The year is the x-value (fixed, known) and the time for the race the y-value.
slope is
change in y/change in x.
It is 3.65-3.80/1980-1960
That is -0.15/20= -0.0075
Now use the point slope formula
y-y1=m(x-x1)
y-3.8=-0.0075(x-1960), using the first point (the second will give the same answer)
y=-.0075x-(-0.0075*1960)+3.8
y=-0.0075x+14.7+3.8
y=-0.0075x+18.5, the first ANSWER
Check using second point.
y-3.65=-0.0075(x-1980)
y-3.65=-0.0075x+14.85
y=-0.0075x+18.5